Integrating factors found by inspection

y(x^2 + y^2 βˆ’ 1)dx + x(x^2 + y^2 + 1)dy = 0 when y(1) = 0


1
Expert's answer
2022-03-24T18:46:03-0400

"y(x^2 + y^2 − 1)dx + x(x^2 + y^2 + 1)dy = 0\\
(x^2y + y^3 − y)dx + (x^3 + xy^2 + x)dy = 0\\
x^2ydx + y^3dx − ydx + x^3dy + xy^2dy + xdy = 0\\
x^2(ydx+xdy)+y^2(ydx+xdy)+xdy-ydx=0\\
\left[(ydx+xdy)(x^2+y^2)+xdy-ydx =0\right]\frac{1}{x^2+y^2}\\
ydx+xdy+ \frac{xdy-ydx}{x^2+y^2}=0\\
d(xy)+d(\arctan (\tfrac{y}{x}))=0\\"


integrating through, we have:

"xy+ \arctan (\tfrac{y}{x})=C"


When y(1) = 0,

"(1)(0) + \arctan (\tfrac{0}{1}) =C\\
0+0=C \implies C=0\\[6mm]
\therefore xy+ \arctan (\tfrac{y}{x})=0"